# Sample Compression Scheme Reductions

### v2

@@ Appendix A Discussion about Infinitized Sample Compression Schemes
A binary tree is perfect if all internal nodes have exactly two children, and all leaf nodes are at the same level. It is said to have depth $d$ if there are $12^{d}-$ 1 vertices. Given a class $\mathcal{C}$, a Littlestone tree is a perfect binary tree, where each vertex is labeled with an element from $\mathcal{X}$. For each path from the root to a leaf, there exists a concept $c\in\mathcal{C}$ such that for each internal node at depth $i$, $c(x_{i})=b_{i}$, where $b_{i}\in\{0,1\}$ indicates whether the path follows the left ($b_{i}=0$) or right ($b_{i}=1$) child. The Littlestone dimension of $\mathcal{C}$ (denoted as $d_{\mathrm{LD}}(\mathcal{C})$) is the maximum depth of a Littlestone tree of $\mathcal{C}$. If no largest depth exists, set $d_{\mathrm{LD}}(\mathcal{C})=\infty$.
@@
Let a binary concept class $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ with a finite Littlestone dimension $d_{\mathrm{LD}}(\mathcal{C})<\infty$. Then, assuming access to the oracle $M(\cdot,\cdot)$, there exists an infinitized sample compression scheme of size $O(d_{\mathrm{LD}}(\mathcal{C}))$ for $\mathcal{C}$.
@@
    • If there exists a fixed $y$ such that $c(x)=y$ for all $c\in\mathcal{C}$, predict $y$.
@@
To prove the compression size, we claim the algorithm will take $d_{\mathrm{LD}}(\mathcal{C})$ iterations. Whenever a mistake is made for some $(x,y)$ at time $t$, we must have that $d_{\mathrm{LD}}(V_{t+1})<d_{\mathrm{LD}}(V_{t})$ ,otherwise, $d_{\mathrm{LD}}(V_{t,x,0})=d_{\mathrm{LD}}(V_{t,x,1})=d_{\mathrm{LD}}(V_{t})=k$ for some $k$, so we can shatter $V_{t}$ using $x$ at the root and the Littlestone trees of depth $k$ for each labeling of $x$, implying a larger Littlestone dimension for $V_{t}$. Thus, the algorithm takes at most $d_{\mathrm{LD}}(\mathcal{C})$ iterations and we have an infinitized compression scheme of size $O(d_{\mathrm{LD}}(\mathcal{C}))$.
@@
Consider the thresholds on the natural numbers. Here $\mathcal{X}=\mathbb{N}$, and $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{X}$ where =$\mathcal{=}\mathopen{}\left\{c_{t}:t\in\mathbb{N}\right\}$ and $c_{t}(x)=\mathbbm{1}[x\leq t]$. This class has an infinite Littlestone dimension. It also has a sample compression scheme of size $2$ – pick the rightmost point with a label equal to 00 and the leftmost point with a label equal to $1$.
@@
Consider the class of thresholds, $\mathcal{C}=\{c_{t}:t\in\mathbb{R}\}$ where $c_{t}(x)=\mathbbm{1}[x\leq t]$. There is no infinitized compression scheme for this.
We show there is no infinitized compression scheme for $\mathcal{C}$. We can assume for now that $\mathcal{C}=\{c_{t}:t\in\mathbb{R}\setminus\mathbb{Q}\}$. Let $S=\mathbb{Q}$. Assume there is an compression scheme ($\kappa,\rho$) for the class where $|\kappa|\leq k$ Notice that for any $t_{1},t_{2}$ irrational, there exists a rational $z$ such that $t_{1}<z<t_{2}$. Thus, $\kappa(S)$ must be different for every $c_{t}\in\mathcal{C}$. Thus, there must be a one-to-one function from $\mathbb{R}\setminus\mathbb{Q}$ to $\cup_{i\leq k}(\{0,1\}\times S)^{i}$. However, the former set is uncountable, while the latter set is countable, so this is impossible.
@@
Consider an infinite domain $\mathcal{X}$, and arrange it as an infinite perfect binary tree. Construct a concept $\mathcal{C}$ as follows. For each $u$ at a finite depth in the tree, consider the path $P_{u}=x_{1},x_{2},\ldots.,x_{k}$. Define the concept $c_{u}$ as follows:
$c_{u}(x)=\begin{cases}1&x\text{ has a right child in $P_{u}$}.\\ 0&\text{ otherwise}.\end{cases}$
@@ Appendix B A Sample Compression Scheme for Classes with Graph Dimension $1$
Ben-David [2015] provides the following sample compression scheme for classes with VC dimension $1$. Consider $S=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ realizable by some $c\in\mathcal{C}$. We can pick a $c_{0}\in\mathcal{C}$ and consider the ordering $\leq$ from Lemma B.3. Identify the deepest point $x_{i}\in\mathopen{}\left\{x\in\mathcal{X}:c(x)\neq c_{0}(x)\right\}$. The compression set is then formed as ${(x_{i},y_{i})}$. For any test point $z$, if there is a unique hypothesis $c\in\mathcal{C}$ consistent with $(x_{i},y_{i})$, i.e., if |$|\mathopen{}\left\{c(z):c(x_{i})=y_{i},c\in\mathcal{C}\right\}|=1$, predict $c(z)$. If multiple hypotheses are consistent, default to predicting according to $c_{0}$.
@@
Consider the following attempt at a compression scheme, given point $(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})$ realizable by some $c\in\mathcal{C}$. Construct the tree from earlier, and compress it to the deepest point $x$ where $g_{c}(x)=1$. For a test point $z$, we can try doing the same as earlier: If there are no ambiguities, predict the only option. Otherwise, predict according to $h_{0}$. However, there are some caveats.
For a test point $z$, if there are no ambiguities, then we are done. Consider the scenario where there are ambiguities. In the binary case, if $x\leq z$, no ambiguities arise as $h(z)\neq h_{0}(z)$. However, for multiclass prediction, additional ambiguities may occur for points $z$ with $x\leq z$. Below, we will provide a way to address this issue.
@@
Input: Concept class $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$, $d_{G}(\mathcal{C})\leq 1$, realizable $S=(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})$.
@@
  • Construct tree ordering $\leq$ from $\{x\in\mathcal{X}:(x,y)\in S\text{ for some }y\in\mathcal{Y}\}$.
@@
  • Let $S_{t}=\{z:x\leq z,|c^{\prime}(z):c^{\prime}\in\mathcal{C}\text{ and }c^{\prime}(x)=c(x)|>1\}$
@@
$\rho(\{x,y\})(z)=\begin{cases}c(z)\text{ for any }c\in\mathcal{C}\text{ consistent with }c(x)=y&\text{ if }|\{c^{\prime}(z):c^{\prime}\in\mathcal{C}\text{ and }c^{\prime}(x)=y\}|=1\\ c_{0}(z)&\text{ otherwise. }\end{cases}$
@@
In the above algorithm, when setting $z_{t+1}$ to be the deepest element in $S_{t}$, $\{c^{\prime}\in\mathcal{C}:c^{\prime}(z_{t+1})=c(z_{t+1})\}\subseteq\{c^{\prime}\in\mathcal{C}:c^{\prime}(z_{t})=c(z_{t})\}$.
@@
Utilizing Lemma B.4, we now proceed with the proof of this lemma. Consider running Algorithm 2. Each time we switch our compression point from $z_{t}$ to $z_{t+1}$, by Lemma B.4, the space of hypotheses consistent with the compression point becomes a strict subset of what it was before. Furthermore, $z_{t+1}\not\in S_{t+1}$, and since $S_{t+1}\supseteq S_{t}$, $S_{t}$ always includes points in $I_{x}$ that are larger than $z_{t}$ in the ordering. Since $I_{x}$ is finite and $S_{t}$ keeps shrinking as we increase $t$, this ends in a finite number of steps until there are no ambiguities when we try to make a prediction for a test point $z\in I_{x}$.
@@
We start with the following Lemma, which will relate the graph dimension and the pseudo-dimension.
@@
For any concept class $\mathcal{C}\subseteq$
$[0,1]^{\mathcal{X}}$, it holds that $d_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$.
@@
Consider $n$ points $S_{n}=\{x_{1},\ldots,x_{n}\}$ that are G-shattered (see Definition 3.1) by $\mathcal{C}$ via values $y_{1},\ldots,\\ y_{n}$, via the $2^{n}$ concepts $\mathcal{C}_{0}\subseteq\mathcal{C}$. Consider $\mathcal{C}_{0}$ supported on $x_{1},\ldots,x_{n}$. We will consider $d_{\mathrm{P}}(\mathcal{C}_{0})$, which will be a lower bound for $d_{\mathrm{P}}(\mathcal{C})$. For any $c\in\mathcal{C}_{0}$, define
@@
$z_{i}=\begin{cases}y_{i}&0\not\in S_{i},\\ y_{i}+\epsilon\textnormal{ for some }\epsilon<\inf_{c\in\mathcal{C}_{n},c(x_{i})>y_{i}}c(x_{i})-y_{i}&1\not\in S_{i},\\ y_{i}-\epsilon\textnormal{ for some }\epsilon<\inf_{c\in\mathcal{C}_{n},c(x_{i})<y_{i}}y_{i}-c(x_{i})&-1\not\in S_{i}.\end{cases}$
Note that $(\mathbbm{1}[c(x_{1})\geq z_{1}],\mathbbm{1}[c(x_{2})\geq y_{2}],\ldots,\mathbbm{1}[c(x_{n})\geq z_{n}])$ is distinct for all $c\in\mathcal{C}_{n}$, so it suffices to bound $|\mathcal{C}_{n}|$ and apply Sauer’s Lemma [Sauer, 1972, Vapnik and Chervonenkis, 1971], as follows: By the construction of $\mathcal{C}_{i}$ for each $i$, since we remove the smallest set in a partition into three at each step, it follows that $|\mathcal{C}_{i}|\geq\frac{2}{3}|\mathcal{C}_{i-1}|$. Thus, $|\mathcal{C}_{n}|\geq 2^{n}(2/3)^{n}$ Let $d_{\mathrm{P}}:=d_{\mathrm{P}}(\mathcal{C}_{n})$. By Sauer’s lemma (taking logs), we have that
@@
The VC dimension of a partial concept class is the largest nonnegative integer $n$ for which there exist $x_{1},x_{2},\ldots,x_{n}$
shattered by $\mathcal{C}.$
@@
Given a class $\mathcal{C}$, a compression scheme is a partial concept class compression scheme if for any sequence $S$ realizable by $\mathcal{C}$, $\rho(\kappa(S))(x)=y$ for all $(x,y)\in S$.
@@
Lemma D.3 (Partial Concept Class with VC $1$ and Unbounded Compression Size [Alon et al., 2022])
@@
In the proof of 5.6, we make use of the following class $\mathcal{C}$ and perturbation set function $\mathcal{U}$: Let each $x\in\mathcal{X}$ have a unique twin $x^{\prime}$ outside of $\mathcal{X}$. Define $\mathcal{X}^{\prime}$ to be $\{x^{\prime}:x\in\mathcal{X}\}$, and $\widetilde{\mathcal{X}}$ to be $\mathcal{X}\cup\mathcal{X}^{\prime}$. Construct $\mathcal{C}$ as follows: $\mathcal{C}=\{g_{c}:c\in\mathcal{C}_{part}\}$ where for $x\in\mathcal{X}$, if $x\in\operatorname*{supp}(c)$, $g_{c}(x)=g_{c}(x^{\prime})=c(x)$, and for $x\not\in\operatorname*{supp}(c)$, $g_{c}(x)=0$ and $g_{c}(x^{\prime})=1$. For $x\in\mathcal{X}$, let $\mathcal{U}(x)=\mathcal{U}(x^{\prime})=\{x,x^{\prime}\}$.
@@
Given a perturbation function $\mathcal{U}$, a class $\mathcal{C}$ is robustly learnable if, for any $\epsilon,\delta>0$ and any robustly realizable distribution $\mathcal{P}$, there exists $n={\mathrm{poly}}(\frac{1}{\delta},\frac{1}{\epsilon})$  [^3] There are many definitions of learnability, and it is standard to let $n={\mathrm{poly}}(d_{VC}(\mathcal{C}),\log\frac{1}{\delta},\frac{1}{\epsilon})$. We adapt the definition here since our concept class has infinite VC dimension, and we relax the $\log\frac{1}{\delta}$ to $\frac{1}{\delta}$, such that, given $n$ i.i.d. samples from $\mathcal{P}$, there is an algorithm that returns a hypothesis $\widehat{c}\in\mathcal{Y}^{\mathcal{X}}$ satisfying
@@
with probability at least $1-\delta$ over the sample.
To show that the class is learnable, we use of the one-inclusion graph predictor, which we define as follows. Note that since we are working in the robust setting, the definition is slightly different, since we require the vertices to correspond to robustly realizable sequences.
@@
Given a concept class $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ and perturbation function $\mathcal{U}$, the One-Inclusion Graph Predictor is an algorithm $\mathcal{A}:(\mathcal{X}\times\mathcal{Y})^{*}\rightarrow\{0,1\}^{\mathcal{X}}$, defined as follows: Given a dataset $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, and test point $z$ consider the following graph: Let the vertices be
$V=\mathopen{}\left\{(c(x_{1}),\ldots,c(x_{n}),c(z)):c\in\mathcal{C},(x_{1},c(x_{1})),(x_{n},c(x_{n})),(z,c(z))\text{ robustly realizable by }\mathcal{C}\right\}.$
@@
$=\mathbbm{E}_{(x_{1},y_{1}),\ldots,(x_{n+1},y_{n+1})\sim\mathcal{P},i\sim\{1,2,\ldots,n+1\}}\bigg{[}\mathbbm{1}[\widehat{c}_{-i}(x_{i})\neq y_{i}]\bigg{]}$
$=\mathbbm{E}_{(x_{1},y_{1}),\ldots,(x_{n+1},y_{n+1})\sim\mathcal{P},i\sim\{1,2,\ldots,n+1\}}\bigg{[}\mathbbm{1}\bigg{[}\bigg{(}(y_{1},..,y_{i},..,y_{n+1}),(y_{1},..,1-y_{i},..,y_{n+1})\bigg{)}\in E(G)\bigg{]}\bigg{]}$

### v3

@@ Appendix A Discussion about Infinitized Sample Compression Schemes
A binary tree is perfect if all internal nodes have exactly two children, and all leaf nodes are at the same level. It is said to have depth $d$ if there are $12^{d}-$ 1 vertices. Given a class $\mathcal{C}$, a Littlestone tree is a perfect binary tree, where each vertex is labeled with an element from $\mathcal{X}$. For each path from the root to a leaf, there exists a concept $c\in\mathcal{C}$ such that for each internal node at depth $i$ with label $x_{i}$, $c(x_{i})=b_{i}$, where $b_{i}\in\{0,1\}$ indicates whether the path follows the left ($b_{i}=0$) or right ($b_{i}=1$) child. The Littlestone dimension of $\mathcal{C}$ (denoted as $d_{\mathrm{LD}}(\mathcal{C})$) is the maximum depth of a Littlestone tree of $\mathcal{C}$. If no largest depth exists, set $d_{\mathrm{LD}}(\mathcal{C})=\infty$.
@@
Let a binary concept
class $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ with a finite Littlestone dimension $d_{\mathrm{LD}}(\mathcal{C})<\infty$. Then, assuming access to the oracle $M(\cdot,\cdot)$, there exists an infinitized sample compression scheme (Algorithm 1) of size $O(d_{\mathrm{LD}}(\mathcal{C}))$ for $\mathcal{C}$.
@@
    • If there exists a fixed $y$ such that $c(x)=y$ for all $c\in\mathcal{C}_{T}$, predict $y$.
@@
To prove the compression size, we claim the algorithm will take $d_{\mathrm{LD}}(\mathcal{C})$ iterations. Whenever a mistake is made for some $(x,y)$ at time $t$, we must have that $d_{\mathrm{LD}}(V_{T_{t+1}})<d_{\mathrm{LD}}(V_{T_{t}})$, since otherwise, $d_{\mathrm{LD}}(V_{T_{t},x,0})=d_{\mathrm{LD}}(V_{T_{t},x,1})=d_{\mathrm{LD}}(V_{T_{t}})=k$ for some $k$, so we can shatter $V_{T_{t}}$ using $x$ at the root and the Littlestone trees of depth $k$ for each labeling of $x$, implying a larger Littlestone dimension for $V_{T_{t}}$. Thus, the algorithm takes at most $d_{\mathrm{LD}}(\mathcal{C})$ iterations and we have an infinitized compression scheme of size $O(d_{\mathrm{LD}}(\mathcal{C}))$.
@@
Consider the thresholds on the natural numbers. Here $\mathcal{X}=\mathbb{N}$, and $\mathcal{C}\subseteq\mathopen{}\left\{0,1\right\}^{\mathcal{X}}$ where $\mathcal{C}=\mathopen{}\left\{c_{t}:t\in\mathbb{N}\right\}$ and $c_{t}(x)=\mathbbm{1}[x\leq t]$. This class has an infinite Littlestone dimension. It also has a sample compression scheme of size $2$ – pick the rightmost point with a label equal to 00 and the leftmost point with a label equal to $1$.
@@
Consider the class of thresholds, $\mathcal{C}=\{c_{t}:t\in\mathbb{R}\}$ where $c_{t}(x)=\mathbbm{1}[x\leq t]$. There is no infinitized compression scheme for this class.
We show there is no infinitized compression scheme for $\mathcal{C}$. We can assume for now that $\mathcal{C}=\{c_{t}:t\in\mathbb{R}\setminus\mathbb{Q}\}$. Let $S=\mathbb{Q}$. Assume there is an compression scheme ($\kappa,\rho$) for the class where $|\kappa|\leq k$. Notice that for any $t_{1},t_{2}$ irrational, there exists a rational $z$ such that $t_{1}<z<t_{2}$. Thus, $\kappa(S)$ must be different for every $c_{t}\in\mathcal{C}$. Thus, there must be a one-to-one function from $\mathbb{R}\setminus\mathbb{Q}$ to $\cup_{i\leq k}(\{0,1\}\times S)^{i}$. However, the former set is uncountable, while the latter set is countable, so this is impossible.
@@
We
construct a concept class $\mathcal{C}$ that has an infinite Littlestone tree (and thus infinite Littlestone dimension) yet admits a simple infinitized compression scheme of size 1.
Consider an infinite domain $\mathcal{X}$ arranged as an infinite perfect binary tree. For each node $u$ at a finite depth in the tree, let $P_{u}=x_{1},x_{2},\ldots,x_{k}$ be the path from the root to $u$. Define the concept $c_{u}$ as follows:
$c_{u}(x)=\begin{cases}1&x\text{ has a right child in $P_{u}$}\\ 0&\text{ otherwise}\end{cases}$
@@ Appendix B A Sample Compression Scheme for Classes with Graph Dimension $1$
Ben-David [2015] provides the following sample compression scheme for classes with VC dimension $1$. Consider $S={(x_{1},y_{1}),\ldots,(x_{n},y_{n})}$ realizable by some $c\in\mathcal{C}$. We can pick a $c_{0}\in\mathcal{C}$ and consider the ordering $\leq$ from Lemma B.3. Identify the deepest point $x_{i}\in\mathopen{}\left\{x\in\mathcal{X}:c(x)\neq c_{0}(x)\right\}$. The compression set consists of the single point $(x_{i},y_{i})$. For any test point $z$, if there exists exactly one possible label value among concepts consistent with $(x_{i},y_{i})$ (that is, if |$|\{c(z):c(x_{i})=y_{i},$
$c\in\mathcal{C}\}|=1$), predict that unique value. Otherwise, default to predicting according to $c_{0}$.
@@
Consider the following attempt at a compression scheme, given points $(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})$ realizable by some $c\in\mathcal{C}$. Construct the tree from earlier (constructed from some $c_{0}\in\mathcal{C}$), and compress to the deepest point $x$ where $g_{c}(x)=1$. For a test point $z$, we can try doing the same as earlier: If there are no ambiguities, predict the only option. Otherwise, predict according to $c_{0}$. However, there are some caveats.
For a test point $z$, if there are no ambiguities, then we are done. Consider the scenario where there are ambiguities. In the binary case, if $x\leq z$, no ambiguities arise as $c(z)\neq c_{0}(z)$. However, for multiclass prediction, additional ambiguities may occur for points $z$ with $x\leq z$. Below, we will provide a way to address this issue.
@@
Input: Concept class $\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$, $d_{G}(\mathcal{C})\leq 1$, realizable $S=(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})$ via some $c\in\mathcal{C}$.
@@
  • Construct tree ordering $\leq$ from $\{x_{1},\ldots,x_{n}\}$ (As in Lemma B.3 statement, from some $c_{0}\in\mathcal{C}$).
@@
  • Let $S_{t}=\{z:z_{t}\leq z,|c^{\prime}(z):c^{\prime}\in\mathcal{C}\text{ and }c^{\prime}(z_{t})=c(z_{t})|>1\}$
@@
$\rho(\{x,y\})(z)=\begin{cases}h(z)\text{ for any }h\in\mathcal{C}\text{ with }h(x)=y&\text{ if }|\{h^{\prime}(z):h^{\prime}\in\mathcal{C}\text{ and }h^{\prime}(x)=y\}|=1\\ c_{0}(z)&\text{ otherwise. }\end{cases}$
@@
In Algorithm 2, when setting $z_{t+1}$ to be the deepest element in $S_{t}$, $\{c^{\prime}\in\mathcal{C}:c^{\prime}(z_{t+1})=c(z_{t+1})\}\subseteq\{c^{\prime}\in\mathcal{C}:c^{\prime}(z_{t})=c(z_{t})\}$.
@@
Utilizing Lemma B.4, we now proceed with the proof of this lemma. Consider running Algorithm 2. Each time we switch our compression point from $z_{t}$ to $z_{t+1}$, by Lemma B.4, the space of hypotheses consistent with the compression point becomes a strict subset of what it was before. Furthermore, $z_{t+1}\not\in S_{t+1}$, and since $S_{t+1}\subseteq S_{t}$, $S_{t}$ keeps shrinking as we increase $t$. All points in $S_{t}$ lie in $I_{z_{1}}$ and are larger than $z_{t}$ in the tree ordering. Since $I_{z_{1}}$ is finite and $S_{t}$ keeps shrinking, this process must end in a finite number of steps when $S_{t}$ becomes empty, at which point there will be no ambiguities in predicting labels for any test point $z\in I_{z_{1}}$.
@@
We start with the following Lemma, which will relate the graph dimension and the
pseudo-dimension.
@@
For any concept class
$\mathcal{C}\subseteq[0,1]^{\mathcal{X}}$, it holds that $d_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$.
@@
Consider $n$ points $S_{n}=\{x_{1},\ldots,x_{n}\}$ that are G-shattered (see Definition 3.1) by $\mathcal{C}$ via values $y_{1},\ldots,\\ y_{n}$, via $2^{n}$ concepts $\mathcal{C}_{0}\subseteq\mathcal{C}$. Consider $\mathcal{C}_{0}$ supported on $x_{1},\ldots,x_{n}$. We will consider $d_{\mathrm{P}}(\mathcal{C}_{0})$, which will be a lower bound for $d_{\mathrm{P}}(\mathcal{C})$. For any $c\in\mathcal{C}_{0}$, define
@@
$z_{i}=\begin{cases}y_{i}&0\not\in S_{i}\text{ or }|S_{i}|=1,\\ y_{i}+\epsilon\textnormal{ for some }\epsilon<\inf_{c\in\mathcal{C}_{n},c(x_{i})>y_{i}}c(x_{i})-y_{i}&-1\not\in S_{i},\\ y_{i}-\epsilon\textnormal{ for some }\epsilon<\inf_{c\in\mathcal{C}_{n},c(x_{i})<y_{i}}y_{i}-c(x_{i})&1\not\in S_{i}.\end{cases}$
Note that $(\mathbbm{1}[c(x_{1})\geq z_{1}],\mathbbm{1}[c(x_{2})\geq z_{2}],\ldots,\mathbbm{1}[c(x_{n})\geq z_{n}])$ is distinct for all $c\in\mathcal{C}_{n}$, so it suffices to bound $|\mathcal{C}_{n}|$ and apply Sauer’s Lemma [Sauer, 1972, Vapnik and Chervonenkis, 1971], as follows: By the construction of $\mathcal{C}_{i}$ for each $i$, since we remove the smallest set in a partition into three at each step, it follows that $|\mathcal{C}_{i}|\geq\frac{2}{3}|\mathcal{C}_{i-1}|$. Thus, $|\mathcal{C}_{n}|\geq 2^{n}(2/3)^{n}$ Let $d_{\mathrm{P}}:=d_{\mathrm{P}}(\mathcal{C}_{n})$. By Sauer’s lemma (taking logs), we have that
@@
The VC dimension of a partial concept class is the largest nonnegative integer $n$ for which there exist $x_{1},x_{2},\ldots,x_{n}\text{ shattered by }\mathcal{C}.$
@@
Given a class $\mathcal{C}$, a compression
scheme is a partial concept class compression scheme if for any sequence $S$ realizable by $\mathcal{C}$, $\rho(\kappa(S))(x)=y$ for all $(x,y)\in S$.
@@
Lemma D.3 (Partial Concept Class with VC $1$ and Unbounded Compression [Alon et al., 2022])
@@
In the proof of Theorem 5.6, we make use of the following class $\mathcal{C}$ and perturbation set function $\mathcal{U}$: Let each $x\in\mathcal{X}$ have a unique twin $x^{\prime}$ outside of $\mathcal{X}$. Define $\mathcal{X}^{\prime}$ to be $\{x^{\prime}:x\in\mathcal{X}\}$, and $\widetilde{\mathcal{X}}$ to be $\mathcal{X}\cup\mathcal{X}^{\prime}$. Construct $\mathcal{C}$ as follows: $\mathcal{C}=\{g_{c}:c\in\mathcal{C}_{part}\}$ where for $x\in\mathcal{X}$, if $x\in\operatorname*{supp}(c)$, $g_{c}(x)=g_{c}(x^{\prime})=c(x)$, and for $x\not\in\operatorname*{supp}(c)$, $g_{c}(x)=0$ and $g_{c}(x^{\prime})=1$. For $x\in\mathcal{X}$, let $\mathcal{U}(x)=\mathcal{U}(x^{\prime})=\{x,x^{\prime}\}$.
@@
Given a perturbation function $\mathcal{U}$, a class $\mathcal{C}$ is robustly learnable if, for any $\epsilon,\delta>0$ and any robustly realizable distribution $\mathcal{P}$, there exists $n={\mathrm{poly}}(\frac{1}{\delta},\frac{1}{\epsilon})$  [^3] There are many definitions of learnability, and it is standard to let $n={\mathrm{poly}}(d_{d_{\mathrm{VC}}}(\mathcal{C}),\log\frac{1}{\delta},\frac{1}{\epsilon})$. We adapt the definition here since our concept class has infinite VC dimension, and we relax the $\log\frac{1}{\delta}$ to $\frac{1}{\delta}$., such that, given $n$ i.i.d. samples from $\mathcal{P}$, there is an algorithm that returns a hypothesis $\widehat{c}\in\mathcal{Y}^{\mathcal{X}}$ satisfying
@@
with probability at least $1-\delta$ over the training sample.
To show that the class is learnable, we use the one-inclusion graph predictor, which we define as follows. Note that since we are in the robust setting, the definition is slightly different, since we require the vertices to correspond to robustly realizable sequences.
@@
Given a concept class
$\mathcal{C}\subseteq\mathcal{Y}^{\mathcal{X}}$ and perturbation function $\mathcal{U}$, the One-Inclusion Graph Predictor is an algorithm $\mathcal{A}:(\mathcal{X}\times\mathcal{Y})^{*}\rightarrow\{0,1\}^{\mathcal{X}}$, defined as follows: Given a dataset $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$, and test point $z$ consider the following graph: Let the vertices $V$ be
$\mathopen{}\left\{(c(x_{1}),\ldots,c(x_{n}),c(z)):c\in\mathcal{C},(x_{1},c(x_{1})),\ldots,(x_{n},c(x_{n})),(z,c(z))\text{ robustly realizable by }\mathcal{C}\right\}.$
@@
$=\mathbbm{E}_{(x_{1},y_{1}),\ldots,(x_{n+1},y_{n+1})\sim\mathcal{P},i\sim[n+1]\}}\bigg{[}\mathbbm{1}[\widehat{c}_{-i}(x_{i})\neq y_{i}]\bigg{]}$
$=\mathbbm{E}_{(x_{1},y_{1}),\ldots,(x_{n+1},y_{n+1})\sim\mathcal{P},i\sim[n+1]}\bigg{[}\mathbbm{1}\bigg{[}\bigg{(}(y_{1},..,y_{i},..,y_{n+1}),(y_{1},..,1-y_{i},..,y_{n+1})\bigg{)}\in E(G)\bigg{]}\bigg{]}$